KittyGirl:

A spinner is divided into sections of equal size, some of which are red, some are blue, and the remaining are green. The probability of the arrow landing on a section colored red is 25%. The probability of the arrow landing on a section colored blue is 10%. What is the probability of the arrow landing on a section colored green? 10% 25% 35% 65%

KittyGirl:

@jhonyy9

HanAkoSolo:

well think about it, the percentage is out of 100

HanAkoSolo:

So think about it this way: 100 - 10 - 26 = ? You are taking parts out of a whole out

HanAkoSolo:

* 100 - 10 - 25 ^ that's what I meant, sorry!

KittyGirl:

lol

KittyGirl:

65?????

HanAkoSolo:

correct!

Study2Learn:

I believe it's 65% because you add 25% and 10% together, then subtract from 100.

Study2Learn:

I believe it's 65% because you add 25% and 10% together, then subtract from 100.

Study2Learn:

Good job kitty!

KittyGirl:

thanks!

Study2Learn:

Explanation: 25/100= 0.25 10/100=0.10 65/100= 0.65

KittyGirl:

thanks guys!

Study2Learn:

No problem